In a college, some courses contribute more towards an overall GPA than other courses. For example, a science class is worth 4 points; History is worth 2 points; Math is worth 3 points; and English is worth 3 points. The values of the grade letters are as follows, A=4, B=3, C=2, D=1, F=0. What is the GPA of a student who made a "c" in Trigonometry, a "B" in American History, an "A" in Botany, and a "B" in Microbiology?
Found 2 solutions by AnlytcPhil, jjnniicc:
Answer by AnlytcPhil(1116)

In a college, some courses contribute more towards an overall GPA than other
courses. For example, a science class is worth 4 points; History is worth 2
points; Math is worth 3 points; and English is worth 3 points. The values of
the grade letters are as follows, A=4, B=3, C=2, D=1, F=0. What is the GPA of
a student who made a "c" in Trigonometry, a "B" in American History, an "A" in
Botany, and a "B" in Microbiology?
'
GPA =
'
The sum of all products (VALUE OF GRADE LETTER)×(NO. OF POINTS COURSE IS
WORTH)
'
divided by
'
The sum of the all NUMBERS OF POINTS COURSES ARE WORTH.
'
C in trig gives the product 3×2 = 6
B in history gives the product 2×3 = 6
A in botany gives the product 4×4 = 16
B in microbiology gives the product = 4×4 = 16
'
Sum of products = 6+6+16+16 = 44
Sum of numbers of points the courses are worth = 3+2+4+4 = 13
'
GPA = 44/13 = 3.384515385 or 3.38 rounded to hundredths.
'
Edwin

Answer by jjnniicc(1)

the above answer is wrong the b in microbiology only gives you 3 points which makes it 3 x 4 and brings the sum of points to 40
which brings the gpa to GPA = 40/13 = 3.0769 or 3.08